Prove that for all multi-indices a and all functions φ in S, x^(a)φ∈S and ∂ ^(a)φ∈S.
1
Expert's answer
2019-10-07T09:20:18-0400

1) "\lim\limits_{|x|\to+\infty}x^u\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(\frac{\partial^{|\alpha|}\varphi}{\partial x^\alpha}\bigr)=\lim\limits_{|x|\to+\infty}x^u\frac{\partial^{|\alpha+\beta|}\varphi}{\partial x^{\alpha+\beta}}". Since "\varphi\in S", we have "\lim\limits_{|x|\to+\infty}x^u\frac{\partial^{|\alpha+\beta|}\varphi}{\partial x^{\alpha+\beta}}=0", so "\frac{\partial^{|\alpha|}\varphi}{\partial x^\alpha}\in S"

2)Prove by induction on "|\beta|" that "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=\sum\limits_{\gamma\le\inf\{\alpha,\beta\}}\sum\limits_{\delta\le\beta}C_{\gamma\delta}x^{\alpha-\gamma}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}" for some "C_{\gamma\delta}" (if "\varepsilon" and "\rho" are multiindices, then "\varepsilon\le\rho\Leftrightarrow\forall i=1,\ldots,n\ \varepsilon(i)\le\rho(i)" ).

a)If "|\beta|=0", then "\gamma=0" and "\delta=0", so "C_{00}=1"

b)Let it is true for all "|\beta|", where "|\beta|\le k".

c)Consider "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)", where "|\beta|=k+1". Then "\beta(s)=\begin{cases}
\beta'(s),&\text{if $s\neq t$}\\
\beta'(s)+1,&\text{if $s=t$}
\end{cases}" for some "t" and "\beta'", where "|\beta'|=k".

So "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=\frac{\partial}{\partial x_t}\left(\frac{\partial^{|\beta'|}}{\partial x^{\beta'}}\bigl(x^\alpha\varphi\bigr)\right)" . By induction hypothesis we have "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=\frac{\partial}{\partial x_t}\left(\sum\limits_{\gamma'\le\inf\{\alpha,\beta'\}}\sum\limits_{\delta'\le\beta'}C_{\gamma'\delta'}x^{\alpha-\gamma'}\frac{\partial^{|\delta'|}\varphi}{\partial x^{\delta'}}\right)="

"\sum\limits_{\gamma'\le\inf\{\alpha,\beta'\}}\sum\limits_{\delta'\le\beta'}C_{\gamma'\delta'}\left(\frac{\partial x^{\alpha-\gamma'}}{\partial x_t}\frac{\partial^{|\delta'|}\varphi}{\partial x^{\delta'}}+x^{\alpha-\gamma'}\frac{\partial^{|\delta'+1|}\varphi}{\partial x^{\delta'}\partial x_t}\right)"

Let "\gamma(s)=\begin{cases}
\gamma'(s),&\text{if $s\neq t$}\\
\gamma'(s)+1,&\text{if $s=t$}
\end{cases}" and "\delta(s)=\begin{cases}
\delta'(s),&\text{if $s\neq t$}\\
\delta'(s)+1,&\text{if $s=t$}
\end{cases}" , then "\gamma\le\beta" and "\delta\le\beta"

We have "\frac{\partial x^{\alpha-\gamma'}}{\partial x_t}=\begin{cases}
(\alpha(t)-\gamma'(t))x^{\alpha-\gamma},&\text{if $\gamma\le\alpha$}\\
0,&\text{if $\gamma\le\alpha$ is false}
\end{cases}" , so "\frac{\partial x^{\alpha-\gamma'}}{\partial x_t}\frac{\partial^{|\delta'|}\varphi}{\partial x^{\delta'}}=0" if "\gamma\le\alpha" is false, and "\frac{\partial x^{\alpha-\gamma'}}{\partial x_t}\frac{\partial^{|\delta'|}\varphi}{\partial x^{\delta'}}=(\alpha(t)-\gamma(t))\frac{\partial x^{\alpha-\gamma}}{\partial x_t}\frac{\partial^{|\delta'|}\varphi}{\partial x^{\delta'}}" if "\gamma\le\alpha" . Also we have "\gamma\le\beta" and "\delta'\le\delta\le\beta".

Next, "x^{\alpha-\gamma'}\frac{\partial^{|\delta'+1|}\varphi}{\partial x^{\delta'}\partial x_t}=x^{\alpha-\gamma'}\frac{\partial^{|\delta|}\varphi}{\partial x^{\delta}}" , where "\gamma'\le\gamma\le\beta", "\gamma'\le\alpha" and "\delta\le\beta" .

So we obtain that "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)" is linear combination of "x^{\alpha-\gamma}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}" , where "\gamma\le\inf\{\alpha,\beta\}" and "\delta\le\beta".

By principle of mathematical induction we have that "\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=\sum\limits_{\gamma\le\inf\{\alpha,\beta\}}\sum\limits_{\delta\le\beta}C_{\gamma\delta}x^{\alpha-\gamma}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}" for some "C_{\gamma\delta}" .

Then "\lim\limits_{|x|\to+\infty}x^u\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=\lim\limits_{|x|\to+\infty}x^u\sum\limits_{\gamma\le\inf\{\alpha,\beta\}}\sum\limits_{\delta\le\beta}C_{\gamma\delta}x^{\alpha-\gamma}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}="

"=\sum\limits_{\gamma\le\inf\{\alpha,\beta\}}\sum\limits_{\delta\le\beta}C_{\gamma\delta}\lim\limits_{|x|\to+\infty}x^{\alpha-\gamma+u}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}"

Since "\varphi\in S", for every "\gamma" and "\delta" from sum we have "\lim\limits_{|x|\to+\infty}x^{\alpha-\gamma+u}\frac{\partial^{|\delta|}\varphi}{\partial x^\delta}=0", so "\lim\limits_{|x|\to+\infty}x^u\frac{\partial^{|\beta|}}{\partial x^\beta}\bigl(x^\alpha\varphi\bigr)=0". That is "x^\alpha\varphi\in S".